Determine relative abundance of isotope

Element X has two isotopes: X - 100 and X 104. If the atomic mass of X is 101 amu, what is the relative abundance of each isotope in nature?
Answer #1

let a = % abundance x-100
let b = % abundance x-104
a + b = 100
a = 100 - b

101 = ( 100-b) + 104 b / 100

10100 = 10000 - 100 b + 104 b

100 = 4 b

b = 25 %
a = 75 %

answered by: Myriel